# Triad: Vision Foundation Model for 3D Magnetic Resonance Imaging

### v1

@@ Abstract
Vision foundation models (VFMs) are pre-trained on extensive image datasets to learn general representations for diverse types of data. These models can subsequently be fine-tuned for specific downstream tasks, significantly boosting performance across a broad range of applications. However, existing vision foundation models that claim to be applicable to various radiology tasks are mostly pre-trained on 3D computed tomography (CT), which benefits from the availability of extensive 3D CT databases. Significant differences between CT and magnetic resonance imaging (MRI) in imaging principles, signal characteristics, and data distribution may hinder their practical performance and versatility in MRI-specific applications. Here, we propose Triad, a vision foundation model for 3D MRI. Triad adopts a widely used autoencoder architecture to learn robust representations from 131,170 3D MRI volumes and uses organ-independent imaging descriptions to constrain the semantic distribution of the visual modality. The above pre-training dataset is called Triad-131K, which is currently the largest 3D MRI pre-training dataset. We evaluate Triad across three tasks, namely, organ/tumor segmentation, organ/cancer classification, and medical image registration, in two data modalities (within-domain and out-of-domain) settings using 25 downstream datasets. By initializing models with Triad’s pre-trained weights, nnUNet-Triad improves segmentation performance by 6.88% compared to nnUNet-Scratch across 17 datasets. Swin-B-Triad achieves a 3.97% improvement over Swin-B-Scratch in classification tasks across five datasets. SwinUNETR-Triad improves by 4.00% compared to SwinUNETR-Scratch in registration tasks across two datasets. Our study demonstrates that pre-training can maximize performance when the data modalities and organs of upstream and downstream tasks are consistent. This work highlights the value of large-scale pre-training techniques for downstream tasks in 3D MRI. By open-sourcing Triad’s weights, code, and data, we aim to enhance the adaptability and reliability of AI solutions for 3D MRI in radiology.
@@ 1 Introduction
Figure 1: Overview of Triad training and evaluation. a. Triad pre-training strategy. Triad implements the reconstruction task based on autoencoders and uses L1 loss for optimization. Imaging descriptions are embedded into vector space to form a distribution, which serves as a supervisory signal to constrain the distribution of visual modalities using Log-ratio loss kim2019deep . The two losses are optimized simultaneously in a multi-task manner. Triad is then evaluated in within-domain tasks and out-of-domain tasks. These include within-domain 3D MRI segmentation, classification, and registration tasks (tasks b, c, and d). And unseen 3D CT/MRI segmentation, classification, and registration tasks (tasks e, f, and g).
@@
Figure 2: An overview of the Triad-131K pre-training dataset. a. Describes the name and scale distribution of each dataset in Triad-131K. b. We compare the parameter scale and data scale used by Triad and existing foundation models, and it is obvious that Triad crushes existing models on both scales. c. Shows examples of visual volumetric modality and textual modality in Triad-131K. d. Shows the dataset scale distribution of three organs: brain, breast, and prostate.
@@
Unfortunately, the potential of foundational models specifically for 3D MRI remains largely unexplored. There are two key limitations: Firstly, although previous general medical foundation models assert the capability to generalize to 3D MRI, substantial differences in imaging principles, signal characteristics, and data distribution between other modalities and MRI may hinder their practical performance and generalizability in MRI-specific applications perez2021torchio ; cardoso2022monai ; https://doi.org/10.1002/mp.17675 . For instance, notable general medical foundation models wu2024voco ; li2024well ; tang2022self have predominantly been pre-trained on 3D computed tomography (CT). Although some foundation models leverage mixed modalities during the pre-training phase—including MRI, CT, positron emission tomography, and microscopy ye2024continual ; gao2024training ; isensee2021nnu —there exists a substantial data imbalance. For example, MedCoSS ye2024continual relied on only 3,000 MRI scans, compared to 10,000 CT scans. Secondly, while there have been attempts to develop 3D MRI-specific foundation models, these efforts have typically focused on data from a single organ cox2024brainsegfounder ; kim2023empirical , primarily emphasizing on T1- (T1-w) and T2-weighted (T2-w) images while overlooking the significance of other MRI sequences. For instance, Brainsegfounder cox2024brainsegfounder uses brain T1-w and T2-w images from approximately 80,000 healthy subjects for pre-training, followed by self-supervised fine-tuning on specific downstream datasets. Moreover, text-based reports blankemeier2024merlin ; zhang2024generalist , which are commonly employed as auxiliary information in 3D CT pre-trained models, are often lacking for 3D MRI data, further limiting the development of comprehensive models. Therefore, the main challenge in building a general 3D MRI vision foundation model is collecting and curating a sufficiently large and diverse dataset. This dataset must cover a wide range of imaging modalities and hardware specifications to ensure the robustness and generalizability of the model.
@@ 2 Results
Figure 3: Study on within-domain 3D tumor segmentation. a. Image segmentation with encoder-decoder architecture by loading the weights of Triad. b. We compare the performance of Scratch, VoCo-SSL and Triad on 5 within-domain datasets based on 3 architectures: Swin-B/L/H. c. We select the nnUNet and Swin-Transformer-Base architectures, along with 3 different weight-loading strategies, and analyze their cross-effects on performance across 5 whitin-domain datasets. d. Consistent with the setting in subfig. c., the radar chart of each category shows the advantage of Triad in tumor segmentation.
@@
We first evaluated the effectiveness of Triad on 3D organ/tumor segmentation, a common tasks in medical image analysis. As shown in Fig. 3 (a), we initialize the encoder with the parameters learned during pre-training, while the decoder is randomly initialized. We evaluate Triad on 17 extensive 3D MRI and CT semantic segmentation datasets, including five MRI datasets for within-domain tasks: BraTS21(baid2021rsna, ), MSD(antonelli2022medical, )-BrainTumour, BreastDM(zhao2023breastdm, ), Prostate158(adams2022prostate158, ), and MSD-Prostate; and 12 datasets covering different organs or modalities for out-of-domain tasks: MM-WHS-MRI(zhuang2018multivariate, ), ATLAS-MRI(quinton2023tumour, ), Abdomen 1K(ma2021abdomenct, ), Kipa22(he2021meta, ), MSD-Pancreas, MSD-Liver, MSD-Heart, MSD-Hippocampus, MSD-Lung, MSD-HepaticVessel, MSD-Spleen, and MSD-Colon. We use the Dice Similarity Coefficient (DSC) as the primary evaluation metric, consistent with public benchmarks.
@@ 2.1.1 Influence of model parameter scale on model performance
It is widely believed that increasing model parameters enhances the performance of downstream tasks in foundation models (oquab2023dinov2, ; chen2024towards, ; ghesu2022contrastive, ; amadou2024echoapex, ). This trend has been observed in various domains, including natural images (oquab2023dinov2, ), X-rays (ghesu2022contrastive, ), and other medical imaging modalities(chen2024towards, ; amadou2024echoapex, ). However, some studies have reported contradictory findings, particularly in 3D CT imaging(wu2024voco, ) and vision language models (shi2025we, ; mei2024bigger, ). In this study, we quantitatively analyze the scaling behavior of Triad pre-training. We conduct a series of experiments with varying model architectures to systematically evaluate their impact on performance.
@@ 2.1.3 Out-of-domain organ/tumor segmentation
We further assess whether the knowledge acquired by Triad during pre-training can be effectively transferred to and applied in new modalities or anatomical organs. To achieve this, we select four MRI datasets from other organs: MSD-Heart, MSD-Liver, MM-WHS-MRI (zhuang2018multivariate, ), and ATLAS-MRI (quinton2023tumour, ). Additionally, we incorporated eight CT datasets: MSD-Hippocampus, MSD-Lung, MSD-Pancreas, MSD-HepaticVessel, MSD-Spleen, MSD-Colon, Abdomen 1K (ma2021abdomenct, ), and Kipa22 (he2021meta, ). We continue to use the nnUNet and SwinUNETR (Swin-B) architectures and employ three parameter initialization methods: training from scratch, VoCo-SSL, and Triad.
@@ 2.2.2 Out-of-domain organ/cancer classification
As illustrated in Fig. 5 (c), Triad achieves the highest performance in the organ classification task and ranks second in both lung nodule and liver lesion classification tasks. Notably, Triad still outperforms training from scratch by +1.02%, demonstrating its effectiveness in generalizing across diverse imaging modalities and organ types. Furthermore, we provide the confusion matrix for Swin-B-Triad across the five datasets. Fig. 5 b (1) shows that when Swin-B-Triad is applied to an out-of-domain classification task with a highly imbalanced category distribution, the model struggles to classify minority classes accurately. In the OrganMNIST3D classification task, Triad fails to distinguish categories 1–4 accurately. These findings suggest that while pre-trained parameters enhance overall downstream performance, addressing challenges such as data imbalance and hard example mining may require specialized sampling strategies or model architectures. Additionally, we present the ROC curves for four datasets in Fig. 5 (d). The ROC curves of all pre-trained models exhibit significant overlap, whereas models trained from scratch show markedly inferior performance, particularly on OrganMNIST3D and ADNI.
@@ 3 Discussion
In this study, we constructed a large-scale 3D MRI dataset, known as TriadMR-131K, which consists of 131,170 volumes from 19,721 patients across 36 clinical datasets. This extensive collection includes a diverse collection of 3D MRI data from three organs, including the brain, breast, and prostate. It includes modalities such as T1-w, T2-w, FLAIR, DWI-MRI, fMRI, and DCE-MRI. Additionally, we extract imaging descriptions from the metadata of each 3D volume. These descriptions detail the imaging modality and associated device parameters. Using this dataset, we develop Triad, a vision foundation model tailored for 3D MRI. Triad employs widely used autoencoder architecture to learn robust representations and incorporates organ-independent imaging descriptions to constrain the semantic distribution of the visual modality. We evaluate Triad on three tasks, including organ and tumor segmentation, organ and cancer classification, and medical image registration. These tasks are assessed across two data modalities—within-domain and out-of-domain— across 25 downstream datasets. By initializing models with Triad’s pre-trained weights, nnUNet-Triad improves segmentation performance by 6.88% over nnUNet-Scratch across 17 datasets. Swin-B-Triad achieves a 3.97% improvement over Swin-B-Scratch in classification tasks across five datasets. SwinUNETR-Triad improves by 4.00% compared to SwinUNETR-Scratch in registration tasks across three datasets. Triad outperforms baseline models across all downstream tasks and exceeds existing state-of-the-art models in most cases. Overall, Triad’s seamless adaptability across downstream tasks highlights its potential as a versatile and efficient tool for diverse clinical applications. This adaptability paves the way for enhanced diagnostic accuracy in 3D MRI.
@@
  • Most vision foundation models for medical image analysis have been pre-trained on large-scale 3D CT datasets (tang2022self, ; ye2024continual, ; zhuang2024mim, ; jiang2023anatomical, ), highlighting the prevalence of CT-based self-supervised learning in this domain. However, our experimental results indicate that pre-training on modality-specific upstream data is more effective. This is demonstrated by the superior performance of Triad, which was pre-trained on MRI data, in 3D MRI-based segmentation, classification, and registration tasks. This finding underscores a critical gap in current research, namely the absence of large-scale 3D MRI pre-training datasets for foundation models. By assembling a diverse collection of 3D MRI data, our work addresses this gap and establishes a more appropriate pre-training paradigm for 3D MRI-based medical imaging tasks. Furthermore, although this study employs an autoencoder (AE) architecture that is not the most advanced in self-supervised learning, our comparison with VoCo-SSL, a state-of-the-art distilled model, suggests that even a relatively simple AE-based framework can outperform more complex models under certain conditions. This highlights an important insight: data quality and task-specific alignment are more critical than model complexity, suggesting that the future of medical AI should emphasize comprehensive and representative pre-training datasets rather than focusing solely on architectural advancements.
@@
Future research can focus on the following aspects. First, improving data quality. Despite having implemented various preprocessing pipelines, the presence of low-quality cases remains a challenge. Numerous studies have emphasized the crucial role of data quality in pre-training (yang2024freemask, ; kirillov2023segment, ). In future work, we plan to design robust automated screening pipelines to screen each volume and enhance data quality. Second, assigning structured electronic health reports to each 3D MRI volume in the Triad-131K dataset to facilitate pre-training for vision-language foundation models. Additionally, expanding downstream tasks to include assistive report generation, visual question answering, and cross-modal medical image retrieval. Lastly, the current pre-training strategy is not limited to MRI scans. Moving forward, we plan to integrate Triad-131K with the largest available CT, PET, X-ray, and ultrasound datasets to pre-train foundation models that generalize effectively across a broad range of radiology tasks, rather than being confined to specific imaging modalities.

### v2

@@ Abstract
Vision foundation models (VFMs) are pre-trained on extensive image datasets to learn general representations for diverse types of data. These models can subsequently be fine-tuned for specific downstream tasks, significantly boosting performance across a broad range of applications. However, existing vision foundation models that claim to be applicable to various clinical tasks are mostly pre-trained on 3D computed tomography (CT), which benefits from the availability of extensive 3D CT databases. Significant differences between CT and magnetic resonance imaging (MRI) in imaging principles, signal characteristics, and data distribution may hinder their practical performance and versatility in MRI-specific applications. Here, we propose Triad, a vision foundation model for 3D MRI. Triad adopts a widely used autoencoder architecture to learn robust representations from 131,170 3D MRI volumes and uses organ-independent imaging descriptions to constrain the semantic distribution of the visual modality. The above pre-training dataset is called Triad-131K, which is currently the largest 3D MRI pre-training dataset. We evaluate Triad across three tasks, namely, organ/tumor segmentation, organ/cancer classification, and medical image registration, in two data modalities (within-domain and out-of-domain) settings using 25 downstream datasets. By initializing models with Triad’s pre-trained weights, nnUNet-Triad improves segmentation performance by 2.51% compared to nnUNet-Scratch across 17 datasets. Swin-B-Triad achieves a 3.97% improvement over Swin-B-Scratch in classification tasks across five datasets. SwinUNETR-Triad improves by 4.00% compared to SwinUNETR-Scratch in registration tasks across two datasets. Our study demonstrates that pre-training can improve performance when the data modalities and organs of upstream and downstream tasks are consistent. This work highlights the value of large-scale pre-training techniques for downstream tasks in 3D MRI. By open-sourcing Triad’s weights, code, and data, we aim to enhance the adaptability and reliability of AI solutions for 3D MRI in clinical tasks.
@@ 1 Introduction
Figure 1: Overview of Triad training and evaluation. a. Triad pre-training strategy. Triad implements the reconstruction task based on autoencoders and uses L1 loss for optimization. Imaging descriptions are embedded into vector space to form a distribution, which serves as a supervisory signal to constrain the distribution of visual modalities using Log-ratio loss kim2019deep . The two losses are optimized simultaneously in a multi-task manner. Triad is then evaluated across within-domain tasks and out-of-domain tasks. These include within-domain 3D MRI segmentation, classification, and registration tasks (tasks b, c, and d). And unseen 3D CT/MRI segmentation, classification, and registration tasks (tasks e, f, and g).
@@
Figure 2: An overview of the Triad-131K pre-training dataset. a. Describes the name and scale distribution of each dataset in Triad-131K. b. We compare the parameter scale and data scale used by Triad and existing foundation models, and it is obvious that Triad surpasses the existing models on both scales. c. Shows examples of visual volumetric modality and textual modality in Triad-131K. d. Shows the dataset scale distribution of three organs: brain, breast, and prostate.
@@
Currently, the potential of foundational models specifically for 3D MRI remains largely unexplored. There are two key limitations: Firstly, although previous general medical foundation models assert the capability to generalize to 3D MRI, substantial differences in imaging principles, signal characteristics, and data distribution between other modalities and MRI may hinder their practical performance and generalizability in MRI-specific applications perez2021torchio ; cardoso2022monai ; https://doi.org/10.1002/mp.17675 . For instance, notable general medical foundation models wu2024voco ; li2024well ; tang2022self have predominantly been pre-trained on 3D computed tomography (CT). Although some foundation models leverage mixed modalities during the pre-training phase, including MRI, CT, positron emission tomography, and microscopy ye2024continual ; gao2024training ; isensee2021nnu , there is extreme data imbalance across the imaging modalities. For example, MedCoSS ye2024continual relied on only 3,000 MRI scans, compared to 10,000 CT scans. Secondly, while there have been attempts to develop 3D MRI-specific foundation models, these efforts have typically focused on data from a single organ cox2024brainsegfounder ; kim2023empirical , primarily emphasizing on T1- (T1-w) and T2-weighted (T2-w) images while overlooking the additional information of other MRI sequences. For instance, Brainsegfounder cox2024brainsegfounder uses brain T1-w and T2-w images from approximately 80,000 healthy subjects for pre-training, followed by self-supervised fine-tuning on specific downstream datasets. Moreover, text-based reports blankemeier2024merlin ; zhang2024generalist , which are commonly employed as auxiliary information in 3D CT pre-trained models, are often lacking for 3D MRI data, further limiting the development of comprehensive models. Therefore, the main challenge in building a general 3D MRI vision foundation model is collecting and curating a sufficiently large and diverse dataset. This dataset must cover a wide range of imaging modalities and hardware specifications to ensure the robustness and generalizability of the model.
@@ 2 Results
Figure 3: Study on within-domain 3D tumor segmentation. a. Image segmentation with encoder-decoder architecture by loading the weights of Triad. b. We compare the performance of Scratch, VoCo-SSL and Triad on 5 within-domain datasets based on 3 architectures: Swin-B/L/H. c. We select the nnUNet and Swin-Transformer-Base architectures, along with 3 different weight-loading strategies, and analyze their cross-effects on performance across 5 within-domain datasets. d. Consistent with the setting in subfig. c., the radar chart of each category shows the overall advantage of Triad in tumor segmentation.
@@
We first evaluated the effectiveness of Triad on 3D organ/tumor segmentation. As shown in Fig. 3 (a), we initialize the encoder with the parameters learned during pre-training, while the decoder is randomly initialized. We evaluate Triad on 17 extensive 3D MRI and CT semantic segmentation datasets, including five MRI datasets for within-domain tasks: BraTS21(baid2021rsna, ), MSD(antonelli2022medical, )-BrainTumour, BreastDM(zhao2023breastdm, ), Prostate158(adams2022prostate158, ), and MSD-Prostate; and 12 datasets covering different organs or modalities for out-of-domain tasks: MM-WHS-MRI(zhuang2018multivariate, ), ATLAS-MRI(quinton2023tumour, ), Abdomen 1K(ma2021abdomenct, ), Kipa22(he2021meta, ), MSD-Pancreas, MSD-Liver, MSD-Heart, MSD-Hippocampus, MSD-Lung, MSD-HepaticVessel, MSD-Spleen, and MSD-Colon. We use the Dice Similarity Coefficient (DSC) as the primary evaluation metric, consistent with public benchmarks.
@@ 2.1.1 Influence of model parameter scale on model performance
It is widely believed that increasing model parameters enhances the performance of downstream tasks in foundation models (oquab2023dinov2, ; chen2024towards, ; ghesu2022contrastive, ; amadou2024echoapex, ). This trend has been observed in various domains, including natural images (oquab2023dinov2, ), x-ray (ghesu2022contrastive, ), and other medical imaging modalities(chen2024towards, ; amadou2024echoapex, ). However, some studies have reported contradictory findings, particularly in 3D CT imaging(wu2024voco, ) and vision language models (shi2025we, ; mei2024bigger, ). In this study, we quantitatively analyze the scaling behavior of Triad pre-training. We conduct a series of experiments with varying model architectures to systematically evaluate their impact on performance.
@@ 2.1.3 Out-of-domain organ/tumor segmentation
We further assess whether the knowledge acquired by Triad during pre-training can be effectively transferred to and applied to other organs or a different imaging modality. To achieve this, we select four MRI datasets from other organs: MSD-Heart, MSD-Liver, MM-WHS-MRI (zhuang2018multivariate, ), and ATLAS-MRI (quinton2023tumour, ). Additionally, we incorporated eight CT datasets: MSD-Hippocampus, MSD-Lung, MSD-Pancreas, MSD-HepaticVessel, MSD-Spleen, MSD-Colon, Abdomen 1K (ma2021abdomenct, ), and Kipa22 (he2021meta, ). We continue to use the nnUNet and SwinUNETR (Swin-B) architectures and employ three parameter initialization methods: training from scratch, VoCo-SSL, and Triad.
@@ 2.2.2 Out-of-domain organ/cancer classification
As illustrated in Fig. 5 (c), Triad achieves the highest performance in the organ classification task and ranks second in both lung nodule and liver lesion classification tasks. Notably, Triad still outperforms training from scratch by +1.02%, demonstrating its effectiveness in generalizing across diverse imaging modalities and organ types. Furthermore, we provide the confusion matrix heydarian2022mlcm for Swin-B-Triad across the five datasets. Fig. 5 b (1) shows that when Swin-B-Triad is applied to an out-of-domain classification task with a highly imbalanced category distribution, the model struggles to classify minority classes accurately. In the OrganMNIST3D classification task, Triad fails to distinguish categories 1–4 accurately. These findings suggest that while pre-trained parameters enhance overall downstream performance, addressing challenges such as data imbalance and hard example mining may require specialized sampling strategies or model architectures. Additionally, we present the ROC curves for four datasets in Fig. 5 (d). The ROC curves of all pre-trained models exhibit significant overlap, whereas models trained from scratch show markedly inferior performance, particularly on OrganMNIST3D and ADNI.
@@ 3 Discussion
In this study, we constructed a large-scale 3D MRI dataset, known as TriadMR-131K, which consists of 131,170 volumes from 19,721 patients across 36 clinical datasets. This extensive collection includes a diverse collection of 3D MRI data from three organs, including the brain, breast, and prostate. It includes modalities such as T1-w, T2-w, FLAIR, DWI-MRI, fMRI, and DCE-MRI. Additionally, we extract imaging descriptions from the metadata of each 3D volume. These descriptions detail the imaging modality and associated device parameters. Using this dataset, we develop Triad, a vision foundation model tailored for 3D MRI. Triad employs widely used autoencoder architecture to learn robust representations and incorporates organ-independent imaging descriptions to constrain the semantic distribution of the visual modality. We evaluate Triad on three tasks, including organ and tumor segmentation, organ and cancer classification, and medical image registration. These tasks are assessed across two data modalities—within-domain and out-of-domain— across 25 downstream datasets. By initializing models with Triad’s pre-trained weights, nnUNet-Triad improves segmentation performance by 2.51% over nnUNet-Scratch across 17 datasets. Swin-B-Triad achieves a 3.97% improvement over Swin-B-Scratch in classification tasks across five datasets. SwinUNETR-Triad improves by 4.00% compared to SwinUNETR-Scratch in registration tasks across two datasets. Triad outperforms baseline models across all downstream tasks and exceeds existing state-of-the-art models in most cases. Overall, Triad’s seamless adaptability across downstream tasks highlights its potential as a versatile and efficient tool for diverse clinical applications. This adaptability paves the way for enhanced diagnostic accuracy in 3D MRI.
@@
  • Most vision foundation models for medical image analysis have been pre-trained on large-scale 3D CT datasets (tang2022self, ; ye2024continual, ; zhuang2024mim, ; jiang2023anatomical, ), highlighting the prevalence of CT-based self-supervised learning in this domain. However, our experimental results indicate that pre-training on modality-specific upstream data is more effective. This is demonstrated by the superior performance of Triad, which was pre-trained on MRI data, in 3D MRI-based segmentation, classification, and registration tasks. This finding underscores a critical gap in current research, namely the absence of large-scale 3D MRI pre-training datasets for foundation models. By assembling a diverse collection of 3D MRI data, our work addresses this gap and establishes a more appropriate pre-training paradigm for 3D MRI-based medical imaging tasks. Furthermore, although this study employs an autoencoder (AE) architecture that is not the most advanced in self-supervised learning, our comparison with VoCo-SSL, a state-of-the-art distilled model, suggests that even a relatively simple AE-based framework can outperform more complex models under certain conditions. This finding highlights an important insight: data quality and task-specific alignment are more critical than model complexity, suggesting that the future of medical AI should emphasize comprehensive and representative pre-training datasets rather than focusing solely on architectural advancements.
@@
Future research can focus on the following aspects. First, improving data quality. Despite having implemented various preprocessing pipelines, the presence of low-quality cases remains a challenge. Numerous studies have emphasized the crucial role of data quality in pre-training (yang2024freemask, ; kirillov2023segment, ). In future work, we plan to design robust automated screening pipelines to screen each volume and enhance data quality. Second, assigning structured electronic health reports to each 3D MRI volume in the Triad-131K dataset to facilitate pre-training for vision-language foundation models. Additionally, expanding downstream tasks to include assistive report generation, visual question answering, and cross-modal medical image retrieval. Lastly, the current pre-training strategy is not limited to MRI scans. Moving forward, we plan to integrate Triad-131K with the largest available CT, PET, X-ray, and ultrasound datasets to pre-train foundation models that generalize effectively across a broad range of clinical tasks, rather than being confined to specific imaging modalities.